[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Aug 29, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Aug 29 12:43:42 UTC 2017


Here is today's summary of NHC computer operations:
--- Pablo

----------

TAFB:
1.) Security cameras remain inoperable at this time. Marc contacted the  
contractor who did not get back to him. William then contacted them again  
late in the day to treat this urgently. They will be stopping by this  
morning.

2) 0345Z Rainfall product did not run for TS Harvey. 0345Z Rainfall product  
for Pot. TC 10 is named "Tropical Storm 10" in the product. Matt Sardi to  
look into this.
HSU:
1.) Despite submitting and retrieving numerous times, can not retrieve EC  
ships for either 93L or 94E at 12Z. Worked for 18z though. Craig and Matt  
Sardi to check this out just in case.

2.) Can we disable the lock screen for the JHT workstation for the national  
TCV/social media desk? This was done already by Brian right away yesterday.  
Thanks Brian.

3.) ATCF running a little slow. Issue with SDM received HWRF request for  
Harvey. Took down all ATCF windows and logged out of all machines around 11  
pm. No further issues reported after that.

4.) Last night ILM notice that in the National TCV NCZ106 was coded with a  
TR.W instead of the TR.A. This is likely due to a discrepancy between that  
zone edit area at NHC and the one at ILM. This is easy to fix but will have  
to wait until after the event. Pablo will follow up on this with Chris  
Mello.


Other items:

NHC RFCs:
See List of active RFCs going on this week send with 8/28 Morning Rounds.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170829/c24fb58a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list